The Information Technology Services (ITS) Alerts System is a rapid-delivery method used to distribute timely information on major ITS services, unexpected outages, and maintenance schedules to the Penn State community.
Updated: 27 min 42 sec ago
Upgrade was completed by 5:45.
Reports of intermittent PASS service degradation are being investigated to determine if an access issue exists with certain PASS gateways. Users may have experienced slow response times when accessing certain gateways over the past few weeks, most notably the win.pass.psu.edu (aka smb.pass.psu.edu, cifs.pass.psu.edu) and nfs.pass.psu.edu gateways.
Staff is working to alleviate the issue that was escalated on December 10.
The small percentage of the VMs that were responding slowly this morning are now functioning at full speed. Below is more information about the degradation that occurred:
When was the degradation resolved?
The degradation was resolved by 12:30 p.m. this afternoon, returning all VMs to normal operating state.
What caused the degradation to occur?
This morning the VM Hosting service team began performing a standard storage migration in response to a storage configuration concern. The migration is routinely done during the day and has never before impacted our users or resulted in any kind of service degradation. As such the migration was evaluated to be a change that could occur outside of our standard maintenance window. However, today, the migration impacted a small percentage of the service\'s VMs, and the service team is continuing to look into why a standard change resulted in an unexpected degradation.
What is being done to prevent a similar degradation from occurring in the future?
Until the root cause of the degradation is identified the VM Hosting service team will be performing any other migrations of this kind during the standard maintenance window.
On December 16, during the maintenance window (5-7 a.m.), an OS update will be made to some UCS servers. Users will experience intermittent service outages as the work is being completed.
No e-mail or information is expected to be lost as a result of the update.
On Wednesday, December 10, 2014 (EST) , there will be a brief outage from 4:00 â 4:30 a.m. (EST) for TrackITS -- "TrackITS Application" & the University Curriculum Archive.
On Thursday, 12/11, beginning at 5:00 AM, the following routers will be reloaded: Blue1, White1, Pollock3 and Redundant3. This is the general deployment of the software that was loaded on Computer3 yesterday. While the new software version did not fully resolve the SNMP software bug, it does have enhanced diagnostics and debug capabilities over the current software, so we will proceed with the general deployment while continuing to work with the vendor on the bug. The routers will be reloaded in sequence following the schedule below:
5:00:01 AM - Pollock3
5:15:01 AM - White1
5:30:01 AM - Red3
5:45:01 AM- Blue1
Turnitin reported a partial service outage on Mon., Dec. 8 from 8:30-9:20 EST. This makes the third outage in December, so we advise instructors to be alert for outages during a deadline. We apologize for the inconvenience.
On Wednesday, December 10, 2014 during the ITS maintenance window (5AM - 7AM), Administrative Information Services (AIS), a unit of Information Technology Services (ITS), will be performing general maintenance on the Web servers that host the AIS Web site (http://ais.its.psu.edu) and Administrative User Guides Web site (http://guides.ais.psu.edu).
During this time, connectivity to the AIS Web site and Administrative User Guides Web site may be degraded or disrupted. All services are expected to be fully restored by 7AM.
Ecommerce is current not accepting payments or transactions.
They are working quickly to resolve the issues.
Sorry for the inconvenience.
Due to the maintenance on the system, eDDS and PeDDS production systems
will not be available between 5am to 7am on Tuesday, December 9, 2014
ITS will be deploying changes for Electronic Account Activation during the maintenance window (5:00 AM - 7:00 AM) on Wednesday, December 10th. A brief service interruption for FPS is expected.
At 5:00 AM on Monday, December 8th, the Computer3 router will be reloaded to pick up a new version of software. This is to address a bug that is affecting SNMP polls. The reload should take less than 5 minutes to complete. If all goes well with the reload and operation of Computer3 for the 48 hour period following the reload, ITS will proceed with a general deployment of that software to the other routers that comprise the new core. (Blue1, White1, Redundant3 & Pollock3) An additional ITS Alert will be posted later regarding the reload of those routers.
FYI - There was a partial Turnitin outage from 5:05-6:05 EST in which users may not have been able to access Originality Reports or GradeMark. Submissions were not affected.
Electronic Theses and Dissertations (eTD) are currently unavailable due to an unexpected outage. Technicians are working on the issue.
iTwo Production (http://itwo.psu.edu) was successfully migrated to the Linux platform. Users can now resume creation, modification, and saving of analyses and dashboards within iTwo.
Please email any questions or issues to firstname.lastname@example.org.
Wednesday morning Dec. 3rd between 5:00AM and 7:00AM, ITS will be performing maintenance work on our ITSP voice circuit. This work has been scheduled with our vendor in effort to facilitate hardware changes. All inbound and outbound voice traffic will route via our redundant ITSP voice circuit.
Service to UCS was restored as of 1:20 p.m. Staff is still looking into the root cause of the issue.
No data was lost as a result of the service outage.
Some users may have experienced problems connecting to Turnitin between 3:20PM and 3:45PM US/Easterb. System is normal now.
The temporary disruption of service was caused by a small database issue, which our Ops team quickly identified and corrected. We apologize for any inconvenience this may have caused.
ITS has been working with our Service Provider to investigate the inbound calling disruption to University Park. Our Service Provider identified and corrected a mis-configured device. As of 1:35 PM, inbound calling capabilities were restored.